OpenStudy (anonymous):

What is the domain of the function f(x)=x+1/2x-1? I absolutely cannot do domain and range for some reason. Please help me!!

OpenStudy (mertsj):

Do you know that you cannot divide by 0?

OpenStudy (anonymous):

yes.

OpenStudy (mertsj):

Do you know, therefore, that the denominator of a fraction cannot be 0?

OpenStudy (anonymous):

yes

OpenStudy (mertsj):

What is the denominator of your fraction?

OpenStudy (anonymous):

2x-1

OpenStudy (mertsj):

So, it would be illegal for 2x-1 to be 0. Do you agree?

OpenStudy (anonymous):

Yes I agree...

OpenStudy (mertsj):

What value of x would therefore be illegal?

OpenStudy (anonymous):

if 2x-1=0 then 2(.5) would equal 1 and 1-1 would equal 0. So x cannot equal .5 right?

OpenStudy (mertsj):

That is correct. So just remember that for a minute.

OpenStudy (mertsj):

Do you know the definition of domain?

OpenStudy (anonymous):

umm... all I know is my teacher usually has us put it in set builder notation. So I would have to say something along the lines of {x|x does not equal 0}

OpenStudy (mertsj):

The domain is the gazillions of values that x CAN be. We know that x CAN'T be 1/2 but it CAN be everything else. And that everything else is the domain.

OpenStudy (mertsj):

So if we start with negative infinity on a number line and go to the right until we get to 1/2...all of those numbers are in the domain. Then we skip 1/2 and go on to the right to positive infinity and all of those numbers are in the domain.

OpenStudy (anonymous):

\[{x|x \in \mathbb{R} \neq .5}\]

OpenStudy (anonymous):

is that what you're saying?

OpenStudy (mertsj):

There are basically three ways to show those numbers: 1. On a number line 2. Interval notation 3. Set builder notation

OpenStudy (mertsj):

And yes, what you wrote is correct is you put the braces around it that make it a set.

OpenStudy (mertsj):

*if

OpenStudy (mertsj):

Do you understand?

OpenStudy (anonymous):

That makes so much more sense... can you be my math teacher from now on? :D

OpenStudy (mertsj):

Sure

OpenStudy (anonymous):

Lol that's amazing! Thanks so much! While you're here... could you help me with another question?

OpenStudy (mertsj):

yes

OpenStudy (anonymous):

Ok so the one I have a question on is: What are the roots of the equation x^2+5x-3=0?

OpenStudy (mertsj):

Well, that won't factor so we have two choices for solving it: 1. Complete the square 2. Use the quadratic formula.

OpenStudy (mertsj):

Which method do you prefer?

OpenStudy (anonymous):

Can we try option 1: complete the square?

OpenStudy (mertsj):

Sure.

OpenStudy (mertsj):

Here are the steps: 1. Move the 3 to the right side.

OpenStudy (mertsj):

2. Divide the coefficient of x by 2

OpenStudy (mertsj):

3. Square that number

OpenStudy (mertsj):

4. Add that number to both sides.

OpenStudy (mertsj):

5. Factor the left side

OpenStudy (mertsj):

6. Take the square root of both sides

OpenStudy (anonymous):

Do you mean divide x^2+5x by 2?

OpenStudy (mertsj):

No. The coefficient of x is 5. Divide 5 by 2

OpenStudy (mertsj):

And just leave it in the form 5/2

OpenStudy (anonymous):

gotcha

OpenStudy (anonymous):

but wouldn't we have to factor the entire side to make it equal and what not? (sorry I'm just making sure I understand you.)

OpenStudy (mertsj):

Yes. After you make it factorable.

OpenStudy (mertsj):

\[x^2+5x=3\] \[x^2+5x+(\frac{5}{2})^2=3+(\frac{5}{2})^2\]

OpenStudy (mertsj):

Now it is factorable

OpenStudy (anonymous):

I see! Thanks! :D

OpenStudy (mertsj):

\[(x+\frac{5}{2})^2=\frac{12}{4}+\frac{25}{4}\]

OpenStudy (mertsj):

\[(x+\frac{5}{2})^2=\frac{37}{4}\]

OpenStudy (mertsj):

Now take the square root of both sides.

OpenStudy (mertsj):

\[\sqrt{(x+\frac{5}{2})^2}=\pm \sqrt{\frac{37}{4}}\]

OpenStudy (mertsj):

\[x+\frac{5}{2}=\pm \frac{\sqrt{37}}{2}\]

OpenStudy (mertsj):

\[x=- \frac{5}{2}\pm \frac{\sqrt{37}}{2}\]

OpenStudy (anonymous):

I get it!!! Awesome! Thank you so much!

OpenStudy (mertsj):

yw